OpenStudy (dls):

Solve: 2tan^-1(cosx)=tan^-1(2cosecx)

OpenStudy (dls):

\[\LARGE 2\tan^{-1}(cosx)=\tan^{-1}(2cosecx)\]

OpenStudy (dls):

can we take tan of both the sides? \[\LARGE \cancel 2(cosx)=\cancel 2(cosecx)\]

hartnn (hartnn):

not at all!

OpenStudy (dls):

why?

hartnn (hartnn):

if you take tan, left = tan (2 arctan(cos x)) which is not 2 cos x

OpenStudy (dls):

tan(arctanx)=x right? and why isn't it 2 cosx ?what is it then?

hartnn (hartnn):

there is obvious difference between tan (arctan) and tan (2 arctan)

OpenStudy (dls):

I was just treating it as a constant,if it was tan2x then things might have been different

hartnn (hartnn):

hmm...constant can't be taken out of inverse trigonometric functions.

OpenStudy (dls):

I see,then how do proceed?

OpenStudy (dls):

to*

hartnn (hartnn):

i can think of this formula : \(\large \arctan x = 2 \arctan \frac{x}{1+\sqrt{1+x^2}}\)

OpenStudy (dls):

haven't heard of it :/

hartnn (hartnn):

now you do, apply it on left side

OpenStudy (dls):

\[\large 2\tan^{-1}x=\tan^{-1} \frac{2x}{1-x^2}\] how about this?

OpenStudy (dls):

\[\large \frac{2cosx}{1-\cos^{2}x}=\frac{2}{sinx}\] gives x=pi/4

hartnn (hartnn):

i have not heard of that formula but ift hats true, yes x= pi/4

OpenStudy (dls):

\[\large \tan2x=\frac{2tanx}{1-\tan^2x}\] you haven't heard of this?!

hartnn (hartnn):

that i know, but what you have written in inverse form , i don't recall